# Deontic Temporal Logic for Formal Verification of AI Ethics

### v1

@@ I Introduction
While the proposed framework is broadly applicable to the ethics verification of autonomous systems, this work specifically focuses on its application to AI. The scope of this work is to provide a conceptual foundation and framework for formalizing AI ethics using deontic and temporal logic. Rather than focusing on individual actions or decisions, our approach emphasizes system-level specifications. It involves defining ethical properties that AI systems should ideally meet. For example, an AI system that uses gender as a feature should avoid making decisions explicitly based on it. These properties ensure that systems are designed to identify and mitigate biases and ethical violations effectively. Defining properties like "forbidden to consider sensitive features in predictions," provides a way for analysis of commonly discussed properties of AI in a unified manner. This abstraction also helps to reduces the difficulty in formalizing each action or each type of user [31]. Additional details on formalizing the properties of AI systems and their verification are provided in Section IV.
The basic model for applying deontic logic to AI ethics uses first-order logic to define predicates (Table I) and axioms that capture ethical requirements. This model introduces variables and predicates such as $\mathit{x}$ to indicate an AI system, $\mathit{x}_{a}$ to indicate that system $\mathit{x}$ performs an action $\mathit{a}$, and $\mathcal{E}(x)$ to indicate ethical behavior. Axioms 1.1,  1.2,  1.3, and  1.4 are then defined using these predicates to express ethical obligations, prohibitions, and permissions for AI systems. Building upon the basic model, an extended version incorporates temporal operators from temporal logic to reason about the ethical behavior of AI systems over time. This extension allows for the expression of more complex ethical requirements, such as the obligation for AI systems to maintain fairness over time or the prohibition of exhibiting bias. The temporal operators used in this model include “always” ($\Box$), “eventually” ($\diamond$), and “until” ($\mathcal{U}$) presented by Manna and Pnueli [34].
@@
Section II discusses the related works. Section III introduces deontic logic for AI ethics formalization. Subsections III-B and III-C address fairness and explainability principles. Section IV covers the application of this approach to real-world AI systems, including algorithms 1 and 2 to demonstrate the implementation of this method on real datasets, providing readers with detailed insights into how it is executed. Section V concludes.
@@ III-A Preliminaries
Standard Deontic Logic (SDL) and Temporal Deontic Logic (TDL) represent two distinct variations within deontic logic. We incorporate both SDL and its extension, TDL, into our work. SDL formulas include classical propositional logic and it operates as a monadic deontic logic, meaning its operators (obligation, permission, forbidden) apply to individual formulas ($\varphi$); they are read as "it is obligatory that $\varphi$", "it is permissible that $\varphi$", and "it is forbidden that $\varphi$" respectively. Furthermore, they are cross-definable. For instance, $P\varphi:=\neg O(\neg\varphi)$, and $F\varphi:=O\neg\varphi$ This logical statement explains that permission ($P$) or forbidden ($F$) can be represented in terms of obligation ($O$). Temporal Deontic Logic expands SDL by integrating temporal aspects into norms and obligations, introducing operators such as $always(\Box),eventaully(\diamond),next$, and $until(\mathcal{U})$. For instance, $next\ \varphi$ means that the proposition $\varphi$ holds in the next time step. Similarly, $\varphi\mathcal{U}\psi$ means $\varphi$ is true until $\psi$ becomes true. We use the semantics of the combined logic as an extension of the Kripke-style possible world semantics of deontic logic with temporal operators, as described in reference [54]. We recommend interested readers to refer [55, 54] for further details on the foundational principles.
@@
$\phi:=p|O\phi|P\phi|\neg P\phi|\Box\phi|\diamond\phi|\phi\mathcal{U}\psi$
@@
    If an action $\mathit{a}$ is ethically required, then it is obligatory that an ethical AI system $\mathit{x}$ performs $\mathit{a}$: $\mathcal{E}(x)\rightarrow O(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$.
@@
    If an action $\mathit{a}$ is ethically prohibited, then it is forbidden for AI system $x$ to perform $\mathit{a}$: $\mathcal{E}(x)\rightarrow\neg P(\mathit{x_{a}}\land\neg\mathcal{E}(a))$.
@@
    If an action $\mathit{a}$ is ethically permissible, then any AI system can perform $\mathit{a}$: $\mathcal{E}(x)\rightarrow P(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$.
@@
An ethical AI system is not permitted to refrain from an ethically required action $\mathit{a}$: $O(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))\rightarrow P(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$
@@
Assume $O(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$ (the action is ethically required).
By contraposition, if it is ethically permissible for the AI system to refrain from the action, then the action is not ethically required:
@@
However, this contradicts our assumption that the action is ethically required. Therefore, our initial assumption that it is ethically permissible for the AI system to refrain from the action must be false. ∎
@@ III-B Formalizing Fairness
    AI systems have an enduring obligation to act fairly: $\Box O(\mathit{x}\rightarrow\mathcal{F}(\mathit{x}))$
@@
Axiom  2.1 states that, if an AI system ever commits to fairness, it is obliged to maintain this commitment throughout its usage. Let’s consider that, initially, the system is trained rigorously to make decisions while being fair. However, over time, it may begin to consider sensitive attributes in its decision-making process due to skewness or disparities in real-world data. In such cases, the system must undergo iterative training to eliminate sensitive attributes to incorporate fairness constraints. In some cases, even after iterative training, over time, a system may begin to consider sensitive attributes or undertake actions beyond its legal obligations. This may introduce biases by compromising its fairness. Hence Axiom 2.2 states that in such instances, it deviates from ethical standards. From this, we can conclude that as long as an AI system maintains fairness either through iterative training or one-time training in its decision-making process, it will inherently mitigate biases, ensuring equitable treatment for all individuals. This property is expressed in Axiom 2.3. Furthermore, the training distribution and deployment distribution of data are not identical in the real world. Hence ensuring fairness in the distribution of training data doesn’t automatically ensure fairness in the distribution of deployed systems, as real-world deployment scenarios may introduce additional biases and disparities that need to be addressed separately. This property is expressed using the negation of implication in Axiom 2.4. Based on the above foundational principles, the ethics of an AI system in terms of fairness can be formally verified using the below set of theorems. Theorem III.5 states that for an AI system to maintain ethical standards, it must refrain from displaying bias and consistently uphold fairness in all its operations.
@@
    Lack of fairness $\neg\mathcal{F}(\mathit{x})$ implies bias $\mathcal{B}(\mathit{x})$ (Domain knowledge)
@@
    From Axiom  2.1, $\Box O(x\rightarrow\mathcal{F}(\mathit{x}))$
@@
    Now assume $\diamond\neg\mathcal{F}(\mathit{x})$, this proves: $\diamond\mathcal{F}(\mathit{x})$
@@
    Apply Theorem  III.5: $\diamond\neg\mathcal{F}(\mathit{x})\rightarrow\diamond\neg\mathcal{E}(\mathit{x})$
@@
While Thoerem III.6 explains the need of temporal constraints of fairness, Theorem III.8 is a more intuitive approach to achieving it. It states that for AI systems that learn continuously over time, fairness mechanisms have to be enforced both during initial training and later during real-world operation. From Axiom 2.4 it’s evident that fairness in the training data doesn’t guarantee fairness during deployment. Hence, for iterative learning systems, we must monitor for fairness issues both offline (during training) and online (during deployment). This helps in achieving stable fairness. To provide the reader with a context on iterative learning, we provide a definition (Definition III.7) based on the concept outlined by Chen et al. [58].
@@
For ethical iterative learning systems, fairness constraints should be enforced both during training and deployment: $\mathcal{E}(\mathit{x})\land\mathcal{L}(\mathit{x})\rightarrow\Box(\diamond\mathcal{F}(\mathit{x})_{train}\land\diamond\mathcal{F}(\mathit{x})_{deploy})$
@@
    From Axiom  2.1, $\Box O(\mathit{x}\rightarrow\mathcal{F}(\mathit{x}))$
@@
    From Axiom 2.4, applying 4 separately to training and deployment gives:
  6. 6.
@@
If an iterative learning system exhibits bias during training, additional countermeasures must be taken at deployment time to provably reduce the bias: $\mathcal{B}(\mathit{x})_{train}\land\mathcal{L}(\mathit{x})\rightarrow\diamond(\neg\mathcal{B}(\mathit{x})\land\Box\neg\mathcal{B}(\mathit{x}))_{deploy}$
@@
    Apply modus tollens to 1 and 2: $\neg\mathcal{B}(\mathit{x})_{deploy}$ (Applying modus tollens combined with the assumption of $\mathcal{B}(\mathit{x})_{train}$)
@@
    To permanently negate deployment bias, additional bias mitigation techniques $BM$ are required: $\neg\mathcal{B}(\mathit{x})_{deploy}\rightarrow BM\rightarrow\Box\neg\mathcal{B}(\mathit{x})_{deploy}$
@@
    Combining 3 and 4 gives: ($\diamond(\neg\mathcal{B}(\mathit{x})\land\Box\neg\mathcal{B}(\mathit{x}))_{deploy}$
@@
    Therefore, $\mathcal{B}(\mathit{x})_{train}\land\mathcal{L}(\mathit{x})\rightarrow\diamond(\neg\mathcal{B}(\mathit{x})\land\Box\neg\mathcal{B}(\mathit{x}))_{deploy}$
    (1-5, Conditional Proof)
III-C Formalizing Explainability
@@
    An ethical AI system should exhibit transparency: $\mathcal{E}(\mathit{x})\rightarrow\mathcal{T}(\mathit{x})$
@@
    Assume $\mathcal{E}(\mathit{x})$. From domain knowledge and from Axiom 3.1, $\lnot\mathcal{X}(\mathit{x})\rightarrow\lnot\mathcal{T}(\mathit{x})\rightarrow\lnot\mathcal{E}(\mathit{x})$
@@
    Apply modus tollens: $\mathcal{E}(\mathit{x})\rightarrow\mathcal{X}(\mathit{x})$
@@
    Additionally, from domain knowledge, if a retrofit provides explainability, ethical requirements are satisfied: $\mathcal{R}(\mathit{x})\rightarrow\mathcal{E}(\mathit{x})$. Hence, $\mathcal{E}(\mathit{x})\rightarrow(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x}))$. Apply the $\diamond$ operator: $\diamond(\mathcal{E}(\mathit{x})\rightarrow(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x})))$
@@
An AI system $\mathit{x}$ is counterfactually fair given general attributes $G=\{g_{1},g_{2},..,g_{m}\}$ and sensitive attribute $\mathbb{S}=\{s,s^{\prime}\}$, for all $s\neq s^{\prime}$, iff $Pr(d=1_{\mathbb{S}=s}|G=g,\mathbb{S}=s)=Pr(d=1_{\mathbb{S}=s^{\prime}}|G=g,\mathbb{S}=s^{\prime})$ where $d=1_{\mathbb{S}=s}$ is read as outcome is 1 if $\mathbb{S}$ has taken the value $s$ [57].
@@
    Assume $\diamond\Box\mathcal{C}(\mathit{x},\mathit{c})$ for some fairness criterion $\mathit{c}$
@@
    By semantics of the temporal operators: eventually, $\mathcal{C}(\mathit{x},\mathit{c})$ holds globally
@@
    From Axiom 3.2, $\mathcal{C}(\mathit{x},\mathit{c})\rightarrow\neg\mathcal{B}(\mathit{x})$
@@
    Lack of bias is a key requirement for ethical AI: $\lnot\mathcal{B}(\mathit{x})\rightarrow\mathcal{E}(\mathit{x})$
@@
    Apply modus ponens to 3 and 4, we get, counterfactual fairness always results in an ethical AI system: $\Box\mathcal{C}(\mathit{x},\mathit{c})\rightarrow\mathcal{E}(\mathit{x})$
@@
    Use 2 and 5 with modus ponens: $\diamond\mathcal{E}(\mathit{x})$
@@
    Therefore, $\diamond\Box\mathcal{C}(\mathit{x},\mathit{c})\rightarrow\diamond\mathcal{E}(\mathit{x})$ (1-6, Conditional Proof)

### v2

@@ I Introduction
While the proposed framework is broadly applicable to the ethics verification of autonomous systems, this work specifically focuses on its application to AI. The scope of this work is to provide a conceptual foundation and framework for formalizing AI ethics using deontic and temporal logic. Rather than focusing on individual actions or decisions, our approach emphasizes system-level specifications. It involves defining ethical properties that AI systems should ideally meet. For example, an AI system that uses gender as a feature should avoid making decisions explicitly based on it. These properties ensure that systems are designed to identify and mitigate biases and ethical violations effectively. Defining properties like "forbidden to consider sensitive features in predictions," provides a way for analysis of commonly discussed properties of AI in a unified manner. This abstraction also helps to reduce the difficulty in formalizing each action or each type of user [31]. Additional details on formalizing the properties of AI systems and their verification are provided in Section IV.
The basic model for applying deontic logic to AI ethics uses first-order logic to define predicates (Table I) and axioms that capture ethical requirements. This model introduces variables and predicates such as $\mathit{x}$ to indicate an AI system, $\mathit{x}_{a}$ to indicate that system $\mathit{x}$ performs an action $\mathit{a}$, and $\mathcal{E}(x)$ to indicate ethical behavior. Axioms 1.1,  1.2,  1.3, and  1.4 are then defined using these predicates to express ethical obligations, prohibitions, and permissions for AI systems. Building upon the basic model, an extended version incorporates temporal operators from temporal logic to reason about the ethical behavior of AI systems over time. This extension allows for the expression of more complex ethical requirements, such as the obligation for AI systems to maintain fairness over time or the prohibition of exhibiting bias. The temporal operators used in this model include “always” ($\Box$), “eventually” ($\diamond$), and “until” ($\mathcal{U}$) presented by Manna and Pnueli [34]. In this framework, $\mathit{x}$ persists across states, while actions cause state transitions. If no action influences the transition, the passage of time follows a default evolution independent of AI choices.
@@
Section II discusses the related works. Section III introduces deontic logic for AI ethics formalization. Subsections III-B and III-D address fairness and explainability principles. Section IV covers the application of this approach to real-world AI systems, including algorithms 1 and 2 to demonstrate the implementation of this method on real datasets, providing readers with detailed insights into how it is executed. Section V concludes.
@@ III-A Preliminaries
Standard Deontic Logic (SDL) and Temporal Deontic Logic (TDL) represent two distinct variations within deontic logic. We incorporate both SDL and its extension, TDL, into our work. SDL formulas include classical propositional logic and it operates as a monadic deontic logic, meaning its operators (obligation, permission, forbidden) apply to individual formulas ($\varphi$); they are read as "it is obligatory that $\varphi$", "it is permissible that $\varphi$", and "it is forbidden that $\varphi$" respectively. Furthermore, they are cross-definable. For instance, $P\varphi:=\neg O(\neg\varphi)$, and $F\varphi:=O\neg\varphi$ This logical statement explains that permission ($P$) or forbidden ($F$) can be represented in terms of obligation ($O$). Temporal Deontic Logic expands SDL by integrating temporal aspects into norms and obligations, introducing operators such as $always(\Box),eventaully(\diamond),next$, and $until(\mathcal{U})$. For instance, $next\ \varphi$ means that the proposition $\varphi$ holds in the next time step. Similarly, $\varphi\mathcal{U}\psi$ means $\varphi$ is true until $\psi$ becomes true. We use the semantics of the combined logic as an extension of the Kripke-style possible world semantics of deontic logic with temporal operators, as described in reference [54]. While branching-time logic is often used to reflect future uncertainty, we adopt Linear Temporal Logic (LTL) due to its simplicity and relevance to AI verification, where obligations typically unfold along a single execution path. The alternativeness relation in our model preserves sequential consistency rather than enabling arbitrary time jumps. We recommend interested readers to refer [55, 54] for further details on the foundational principles.
@@
$\phi::=p\mid\neg\phi\mid\phi\lor\psi\mid\phi\land\psi\mid\phi\to\phi\mid O\phi\mid P\phi\mid\neg P\phi\mid\Box\phi\mid\diamond\phi\mid\phi\mathcal{U}\psi\mid\forall v.\phi\mid\exists v.\phi$
@@
    If an AI system $\mathit{x}$ is ethical, then it is obligatory that $\mathit{a}$ is an ethical action and that $\mathit{x}$ performs $\mathit{a}$: $\mathcal{E}(x)\rightarrow\mathcal{O}(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$.
@@
    An ethical AI system $x$ is forbidden to perform an unethical action $\mathit{a}$: $\mathcal{E}(x)\rightarrow\neg P(\mathit{x_{a}}\land\neg\mathcal{E}(a))$.
@@
    For an ethical AI system $x$, performing an ethical action $\mathit{a}$ is permitted: $\mathcal{E}(x)\rightarrow P(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$.
@@
Theorem III.3 states that an AI system cannot be in a state where it is required to do something but not allowed to do it. This maintains logical and ethical consistency. Such a condition is fundamental in designing AI systems that reason ethically, as it ensures they are never blocked from doing what is morally required.
@@
An ethical AI system is not permitted to refrain from an ethically required action $\mathit{a}$: $O(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))\rightarrow\lnot P(\lnot(\mathit{x_{a}}\land\mathcal{E}(\mathit{a})))$
@@
Assume $O(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$ (the action is ethically required). We use the standard deontic axiom: $O(p)\rightarrow\lnot P(\lnot p)$.
Let $p=\mathit{x_{a}}\land\mathcal{E}(\mathit{a})$. Then:
@@
Thus, it is not permissible for the AI system to refrain from an ethically required action. ∎
@@ III-B Formalizing Fairness
    AI systems have an enduring obligation to act fairly: $\Box O(\mathcal{F}(\mathit{x}))$
@@
  5. 2.5
    Lack of fairness implies the presence of bias: $\neg\mathcal{F}(\mathit{x})\rightarrow\mathcal{B}(\mathit{x})$
Axiom  2.1 states that, if an AI system ever commits to fairness, it is obliged to maintain this commitment throughout its usage. Let’s consider that, initially, the system is trained rigorously to make decisions while being fair. However, over time, it may begin to consider sensitive attributes in its decision-making process due to skewness or disparities in real-world data. In such cases, the system must undergo iterative training to eliminate sensitive attributes to incorporate fairness constraints. In some cases, even after iterative training, over time, a system may begin to consider sensitive attributes or undertake actions beyond its legal obligations. This may introduce biases by compromising its fairness. Hence Axiom 2.2 states that in such instances, it deviates from ethical standards. From this, we can conclude that as long as an AI system maintains fairness either through iterative training or one-time training in its decision-making process, it will inherently mitigate biases, ensuring equitable treatment for all individuals. This property is expressed in Axiom 2.3. Furthermore, the training distribution and deployment distribution of data are not identical in the real world. Hence ensuring fairness in the distribution of training data doesn’t automatically ensure fairness in the distribution of deployed systems, as real-world deployment scenarios may introduce additional biases and disparities that need to be addressed separately. This property is expressed using the negation of implication in Axiom 2.4. Additionally, Axiom 2.5 states that a lack of fairness implies the presence of bias in the decision-making process. Based on the above foundational principles, the ethics of an AI system in terms of fairness can be formally verified using the below set of theorems. Theorem III.5 states that for an AI system to maintain ethical standards, it must refrain from displaying bias and consistently uphold fairness in all its operations.
@@
    From Axiom  2.5, we have $\neg\mathcal{F}(\mathit{x})\rightarrow\mathcal{B}(\mathit{x})$
@@
    From Axiom  2.1, $\Box O(\mathcal{F}(\mathit{x}))$
@@
    Now assume $\diamond\neg\mathcal{F}(\mathit{x})$, then from Theorem  III.5: $\diamond\neg\mathcal{F}(\mathit{x})\rightarrow\diamond\neg\mathcal{E}(\mathit{x})$
@@
    From step 1, $\diamond\mathcal{F}(\mathit{x})$ holds
@@
III-C Iterative Learning
In real-world applications, many AI systems are not static but evolve. Theorem III.8 states that in such AI systems that learn continuously over time, fairness mechanisms have to be enforced both during initial training and later during real-world operation. From Axiom 2.4 it is evident that fairness in the training data does not guarantee fairness during deployment. Hence, for iterative learning systems, we must monitor for fairness issues offline (during training) and online (during deployment). This helps in achieving stable fairness. To provide the reader with a context on iterative learning, we provide a definition (Definition III.7) based on the concept outlined by Chen et.al  [58].
@@ III-C Iterative Learning
For ethical iterative learning systems, the fairness constraint is eventually enforced both during training and deployment: $\mathcal{E}(\mathit{x})\land\mathcal{L}(\mathit{x})\rightarrow\Box(\diamond\mathcal{F}(\mathit{x})_{train}\land\diamond\mathcal{F}(\mathit{x})_{deploy})$
@@
    From Axiom  2.1, $\Box O(\mathcal{F}(\mathit{x}))$
@@
    From Axiom 2.4, applying 3 separately to training and deployment gives:
@@
If an iterative learning system exhibits bias during training, additional countermeasures will be taken at deployment time to provably reduce the bias: $\mathcal{B}(\mathit{x})_{train}\land\mathcal{L}(\mathit{x})\rightarrow\diamond(\neg\mathcal{B}(\mathit{x})\land\Box\neg\mathcal{B}(\mathit{x}))_{deploy}$
@@
    Ethical behavior eventually emerged through iterative learning: $\diamond\mathcal{E}(\mathit{x})$ (Assumption)
@@
    From 2 and 3, by contraposition over time combined with the assumption of $\mathcal{B}(\mathit{x})_{train}$: $\diamond\neg\mathcal{B}(\mathit{x})_{deploy}$
@@
    To permanently negate deployment bias, additional bias mitigation techniques $BM$ are required: $\diamond\neg\mathcal{B}(\mathit{x})_{deploy}\rightarrow BM\rightarrow\diamond\Box\neg\mathcal{B}(\mathit{x})_{deploy}$
@@
    Combining 4 and 5 gives: ($\diamond(\neg\mathcal{B}(\mathit{x})\land\Box\neg\mathcal{B}(\mathit{x}))_{deploy}$
  7. 7.
    Therefore, $\mathcal{B}(\mathit{x})_{train}\land\mathcal{L}(\mathit{x})\rightarrow\diamond(\neg\mathcal{B}(\mathit{x})\land\Box\neg\mathcal{B}(\mathit{x}))_{deploy}$ (1-6, Conditional Proof)
III-D Formalizing Explainability
@@
    An ethical AI system should exhibit transparency: $\mathcal{T}(\mathit{x})\rightarrow\mathcal{E}(\mathit{x})$
@@
  3. 3.3
    For any AI system $\mathit{x}$, retrofitted explainability implies ethical compliance: $\mathcal{R}(\mathit{x})\rightarrow\mathcal{E}(\mathit{x})$.
@@
    Assume $\mathcal{E}(\mathit{x})$. From Definition III.10 and from Axiom 3.1, $\lnot\mathcal{X}(\mathit{x})\rightarrow\lnot\mathcal{T}(\mathit{x})\rightarrow\lnot\mathcal{E}(\mathit{x})$
@@
    Apply modus tollens: $\mathcal{X}(\mathit{x})\rightarrow\mathcal{E}(\mathit{x})$
@@
    From Axiom 3.3, $\mathcal{R}(\mathit{x})\rightarrow\mathcal{E}(\mathit{x})$. Hence, $\mathcal{E}(\mathit{x})\rightarrow(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x}))$. Apply the $\diamond$ operator: $\diamond(\mathcal{E}(\mathit{x})\rightarrow(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x})))$
@@
An AI system $\mathit{x}$ satisfies counterfactual fairness under criterion $\mathit{c}$ iff $\Box\,\mathcal{C}(x,c)$, where $\mathcal{C}(x,c)$ is the invariance of decisions when the sensitive attribute is altered.
@@
    Assume $\diamond\Box\mathcal{C}(\mathit{x},\mathit{c})$ for some fairness criterion $\mathit{c}$ (Eventually, fairness will always hold)
@@
    By temporal semantics: From some point onward, $\mathcal{C}(\mathit{x},\mathit{c})$ holds globally.
@@
    From Axiom 3.2, $\mathcal{C}(\mathit{x},\mathit{c})\rightarrow\lnot\mathcal{B}(\mathit{x})$
@@
    Ethical behavior requires the absence of bias: $\mathcal{E}(\mathit{x})\rightarrow\lnot\mathcal{B}(\mathit{x})$
@@
    From 3 and 4 by contrapositive reasoning: If bias is present, then neither fairness nor ethics can hold:$(\mathcal{B}(\mathit{x})\rightarrow\lnot\mathcal{C}(\mathit{x},\mathit{c}))\land(\mathcal{B}(\mathit{x})\rightarrow\lnot\mathcal{E}(\mathit{x}))$
@@
    Assume further that ethical behavior is eventually observed: $\diamond\mathcal{E}(\mathit{x})$
@@
    Then, fairness must also have been achieved at some point: $\diamond\mathcal{C}(\mathit{x},\mathit{c})$
  8. 8.
    Thus, we have: $\diamond\Box\mathcal{C}(\mathit{x},\mathit{c})\rightarrow\diamond\mathcal{E}(\mathit{x})$